# Mapping 1,000+ Language Models via the Log-Likelihood Vector

### v1

@@ Abstract
To compare autoregressive language models at scale, we propose using log-likelihood vectors computed on a predefined text set as model features. This approach has a solid theoretical basis: when treated as model coordinates, their squared Euclidean distance approximates the Kullback–Leibler divergence of text-generation probabilities. Our method is highly scalable, with computational cost growing linearly in both the number of models and text samples, and is easy to implement as the required features are derived from cross-entropy loss. Applying this method to over 1,000 language models, we constructed a “model map,” providing a new perspective on large-scale model analysis.
@@
Momose Oyama1,2 Hiroaki Yamagiwa1 Yusuke Takase1  Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN {oyama.momose, hiroaki.yamagiwa, y.takase}@sys.i.kyoto-u.ac.jp, shimo@i.kyoto-u.ac.jp
@@ 1 Introduction
meta-llama/Meta-Llama-3-8B KL (/byte) google/codegemma-2b KL (/byte) deepseek-ai/deepseek-llm-7b-base KL (/byte) Undi95/Meta-Llama-3-8B-hf 0.003 (0.000) deepseek-ai/deepseek-coder-1.3b-instruct 1654.767 (2.455) deepseek-ai/deepseek-moe-16b-base 130.715 (0.194) dfurman/Llama-3-8B-Orpo-v0.1 6.994 (0.010) bigcode/starcoderbase-1b 1812.768 (2.690) deepseek-ai/deepseek-llm-7b-chat 245.408 (0.364) migtissera/Tess-2.0-Llama-3-8B 28.829 (0.043) deepseek-ai/deepseek-coder-1.3b-base 2115.462 (3.139) deepseek-ai/deepseek-moe-16b-chat 247.889 (0.368) freewheelin/free-llama3-dpo-v0.2 67.783 (0.101) bigcode/gpt_bigcode-santacoder 2644.348 (3.924) deepseek-ai/DeepSeek-V2-Lite 306.530 (0.455) jondurbin/bagel-8b-v1.0 110.365 (0.164) deepseek-ai/deepseek-coder-6.7b-instruct 2675.004 (3.969) deepseek-ai/ESFT-vanilla-lite 307.487 (0.456) migtissera/Llama-3-8B-Synthia-v3.5 128.141 (0.190) Qwen/CodeQwen1.5-7B-Chat 2759.243 (4.094) deepseek-ai/DeepSeek-V2-Lite-Chat 585.090 (0.868) nvidia/Llama3-ChatQA-1.5-8B 128.776 (0.191) NTQAI/Nxcode-CQ-7B-orpo 2767.578 (4.106) mistralai/Mistral-7B-Instruct-v0.1 913.708 (1.356) ruslanmv/Medical-Llama3-8B 138.566 (0.206) Salesforce/codegen-6B-multi 2859.701 (4.243) statking/zephyr-7b-sft-full-orpo 1089.127 (1.616) FairMind/Llama-3-8B-4bit-UltraChat-Ita 199.748 (0.296) bigcode/starcoderbase-7b 2990.326 (4.437) Severian/ANIMA-Phi-Neptune-Mistral-7B 1140.161 (1.692) NousResearch/Hermes-2-Theta-Llama-3-8B 222.357 (0.330) deepseek-ai/deepseek-coder-6.7b-base 3282.352 (4.870) sethuiyer/Medichat-Llama3-8B 1157.599 (1.718)
Table 1: Top 10 nearest neighbors among the 1,018 language models for each model listed in the first row. The KL divergence is computed using formula (3) in Section 2. The values in parentheses represent the KL divergence measured in bits per byte of text (Section 3.5), obtained by multiplying the KL divergence by 0.001484. Tables for the nearest neighbors of the models labeled in the top panel of Fig. 1 are provided in Appendix I.
@@ 2.2 Log-likelihood vector
For a model $p_{i}$, the probability of generating a text $x$ is denoted $p_{i}(x)$. Following the convention in statistical model selection, we refer to $p_{i}(x)$ as the likelihood of model $p_{i}$ given the text $x$. The log-likelihood is then defined as
@@ 2.4 Kullback-Leibler divergence
The main results, namely (2) and (3), are proved in Appendix D using the theory of exponential family of distributions (Barndorff-Nielsen, 2014; Efron, 1978, 2022; Amari, 1982), similar to the discussion on the relationship between the norm of embeddings and KL divergence (Oyama et al., 2023). Although the concepts of model map and model coordinates have been discussed in statistics (Shimodaira, 1993; Shimodaira and Cao, 1998; Shimodaira, 2001), and there have been a few applications of model maps (Shimodaira and Hasegawa, 2005; Shimodaira and Terada, 2019), they have seldom been utilized in practice.
@@ 2.5 Model coordinates
We primarily use $\bm{q}_{i}$ as the feature vector of model $p_{i}$ and refer to it as the model coordinates. As shown in (3), the squared Euclidean distance in the $\bm{q}$-coordinate system approximates the KL divergence between language models, [^3] That is, $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}\approx 2N\,\mathrm{KL}(p_{i},p_{j})$. For simplicity, we omit the constant scaling factor in expressions of this type. indicating that $\bm{q}_{i}$ represents the position of $p_{i}$ in the space of probability distributions. Since $\bm{\xi}_{i}$ differs from $\bm{q}_{i}$ only by an offset from the origin, $\bm{\xi}_{i}$ also serves as a model coordinate, and $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\|\bm{\xi}_{i}-\bm{\xi}_{j}\|^{2}$. However, we prefer $\bm{q}_{i}$ for its more interpretable components and thus adopt it throughout this paper.
@@ 3 Experimental Setup
Figure 4: Hierarchical clustering of the top 100 most-downloaded models, based on their feature vectors $\bm{q}_{i}$. Model names are color-coded by model type.
@@ 3.1 Selection of text data
The texts used for computing the language models’ coordinates were extracted from the Pile Gao et al. (2020), with five categories of copyrighted material removed. [^4] https://huggingface.co/datasets/monology/pile-uncopyrighted This yielded a dataset $D$ consisting of 10,000 texts, each tagged with a category label from the Pile. Figure 2 visualizes these texts.
@@ 3.2 Selection of language models
We used $K=1{,}018$ language models in total. Of these, 1,000 were selected from models listed on Open LLM Leaderboard v1. Specifically, we considered CausalLM models ranging from 1B to 13B parameters and ranked them by their number of downloads over the 30 days preceding February 1, 2025.555Hugging Face’s API provides the total number of downloads in the past 30 days. We initially selected the top 1,100 models by download count and attempted log-likelihood calculations. Among these, 1,011 successfully produced valid log-likelihood values, and we chose the 1,000 most frequently downloaded from that set. In addition, we included 18 models from the DeepSeek language model series. We obtained information on model parameter sizes and architectures from the Leaderboard. Appendix G provides basic information on the selected models and details on how model types were defined. A complete list of models used in this study is given in Appendix L.
@@ 3.4 Obtaining the leaderboard scores
We obtained benchmark scores for the language models used in our experiments from Open LLM Leaderboard v1.666https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard Between April 2023 and June 2024, this leaderboard evaluated language models on six tasks: AI2 Reasoning Challenge (ARC) Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021a), TruthfulQA Lin et al. (2022a), Winogrande Sakaguchi et al. (2019), and GSM8K Cobbe et al. (2021). Along with individual task scores, we also use the average score across all six tasks, referred to as 6-TaskMean.
@@ 3.5 Byte-normalized KL divergence for cross-experiment comparison
Figure 6: A model map color-coded by (Left) model size and (Right) model creation date.
@@ 4.2 Detection of data leakage
Pearson’s $r$ 0.972 0.948 0.963 0.958 0.965 0.935 0.976 0.994
Spearman’s $\rho$ 0.976 0.974 0.969 0.937 0.973 0.899 0.978 0.990
Table 2: Results of ridge regression for predicting benchmark scores from model coordinates. Predictions for 6-TaskMean and mean log-likelihood are also included. High correlation coefficients are observed across all settings.
@@ 5.1 Benchmark scores and models
We use six benchmark scores from Open LLM Leaderboard v1, as described in Section 3.4. Our experiments are conducted on 996 models for which these benchmark scores are available101010 From the 1,018 models, we excluded four that lacked TruthfulQA scores and 18 from deepseek-ai, as their benchmark scores were incomplete due to the models being relatively new, leaving 996 for analysis. For simplicity, we denote the number of models as $K$. .
@@ 5.2 Setting for regression analysis
For each benchmark task, the dataset is given as $\{(\bm{q}_{1},f_{1}),\dots,(\bm{q}_{K},f_{K})\}$, where $\bm{q}_{i}\in\mathbb{R}^{N}$ is the double-centered log-likelihood vector of the language model $p_{i}$, and $f_{i}\in[0,100]$ is its corresponding benchmark score. We use ridge regression to predict each benchmark score. Let $\bm{Q}\in\mathbb{R}^{K\times N}$ be the matrix of explanatory variables, and let $\bm{f}=(f_{1},\ldots,f_{K})\in\mathbb{R}^{K}$ be the vector for the target variable. The objective function with parameter $\bm{w}\in\mathbb{R}^{N}$ is given by:
$L(\bm{w})=\|\bm{f}-\bm{Q}\bm{w}\|^{2}+\alpha\|\bm{w}\|^{2},$
@@
We split the set of models into five folds and perform parameter training [^11] We used scikit-learn Varoquaux et al. (2015) RidgeCV. and benchmark score prediction. To mitigate the effect of randomness, we repeat the data splitting with five different seeds and take the average of the predictions as the final predicted score. As evaluation metrics, we compute Pearson’s $r$ and Spearman’s $\rho$ to measure the correlation between the predicted and benchmark scores. Additionally, we conduct experiments by replacing the target variable with 6-TaskMean and mean log-likelihood, leading to a total of eight experimental settings. See Appendix K.1 for details.
5.3 Results
Table 2 presents the results of the regression analysis. For all benchmark tasks, the correlation coefficients are high, demonstrating that ridge regression achieves strong predictive performance. Similarly, for 6-TaskMean, the correlation remains high, as also illustrated by the scatter plot in Fig. 8. Even for mean log-likelihood, the regression model attains a high correlation, despite the explanatory variables being double-centered.
Figure 8: Scatter plot of predicted scores versus 6-TaskMean scores for test sets. Pearson’s $r$ is 0.976, and Spearman’s $\rho$ is 0.978. Each point is color-coded by the mean log-likelihood, and the color bar is clipped at the 10th percentile. Higher mean log-likelihood values generally correspond to higher 6-TaskMean scores. However, some models with extremely high mean log-likelihood due to data leakage (Section 4.2) appear in the lower score region, deviating from the general trend. Despite this, the regression-based predictions remain highly accurate, as reflected in the strong correlation coefficients. Scatter plots for individual benchmark tasks are provided in Fig. 18 in Appendix K.2.
@@ 6 Empirical Validation of Theory
Figure 10: Visualization of 36 language models obtained by linearly interpolating pretrained model weights. Each point is color-coded according to its mean log-likelihood. (Left) Models in the weight parameter space. (Right) Models in the probability distribution space, represented by the $\bm{q}$-coordinate system.
In Section 2, we discussed model coordinates for text probability models. Appendix E extends this framework to token-sequence probability models. These theoretical results state that the squared Euclidean distance in the model-coordinate space approximates the KL divergence between models. To validate this, we first evaluate token-level conditional probability models (Section 6.1), since token-level experiments are generally easier to conduct than text-level experiments. We then extend our analysis to text probability models (Section 6.2). Additionally, in Section 6.3, we explore the relationship between model weight parameters and model coordinates.
@@ 6.3 Relationship between model weights and model coordinates
For language models with the same architecture, comparison can also be conducted via weight parameters. We investigated how the structure of the model-coordinate space aligns with the structure of the weight-parameter space. We generated 36 new language models by linearly interpolating the weights of Llama-2-7b-hf, Llama-2-7b-chat-hf (Touvron et al., 2023b), and vicuna-7b-v1.5 (Zheng et al., 2023). For these 36 models, we computed text-generation log-likelihoods and visualized the resulting model coordinates in two dimensions using Principal Component Analysis (PCA). Figure 10 shows these 36 models in both weight space and model coordinate space. We observe that the two-dimensional grid structure in the weight space is mapped continuously into the model coordinate space. This finding supports the validity of (6) and (14) in Appendix D, especially for models that are close to one another. Further details on this experiment can be found in Appendix J.

### v2

@@ Abstract
To compare autoregressive language models at scale, we propose using log-likelihood vectors computed on a predefined text set as model features. This approach has a solid theoretical basis: when treated as model coordinates, their squared Euclidean distance approximates the Kullback-Leibler divergence of text-generation probabilities. Our method is highly scalable, with computational cost growing linearly in both the number of models and text samples, and is easy to implement as the required features are derived from cross-entropy loss. Applying this method to over 1,000 language models, we constructed a “model map,” providing a new perspective on large-scale model analysis. Our code and data are available at https://github.com/shimo-lab/modelmap.
@@
Momose Oyama1,2 Hiroaki Yamagiwa1 Yusuke Takase1  Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN oyama.momose@sys.i.kyoto-u.ac.jp, h.yamagiwa@i.kyoto-u.ac.jp, y.takase@sys.i.kyoto-u.ac.jp, shimo@i.kyoto-u.ac.jp
@@ 1 Introduction
meta-llama/Meta-Llama-3-8B KL [bpb] google/codegemma-2b KL [bpb] deepseek-ai/deepseek-llm-7b-base KL [bpb] Undi95/Meta-Llama-3-8B-hf 4.56e-06 deepseek-ai/deepseek-coder-1.3b-instruct 2.46 deepseek-ai/deepseek-moe-16b-base 0.194 dfurman/Llama-3-8B-Orpo-v0.1 0.0104 bigcode/starcoderbase-1b 2.69 deepseek-ai/deepseek-llm-7b-chat 0.364 migtissera/Tess-2.0-Llama-3-8B 0.0428 deepseek-ai/deepseek-coder-1.3b-base 3.14 deepseek-ai/deepseek-moe-16b-chat 0.368 freewheelin/free-llama3-dpo-v0.2 0.101 bigcode/gpt_bigcode-santacoder 3.92 deepseek-ai/DeepSeek-V2-Lite 0.455 jondurbin/bagel-8b-v1.0 0.164 deepseek-ai/deepseek-coder-6.7b-instruct 3.97 deepseek-ai/ESFT-vanilla-lite 0.456 migtissera/Llama-3-8B-Synthia-v3.5 0.190 Qwen/CodeQwen1.5-7B-Chat 4.09 deepseek-ai/DeepSeek-V2-Lite-Chat 0.868 nvidia/Llama3-ChatQA-1.5-8B 0.191 NTQAI/Nxcode-CQ-7B-orpo 4.11 mistralai/Mistral-7B-Instruct-v0.1 1.356 ruslanmv/Medical-Llama3-8B 0.206 Salesforce/codegen-6B-multi 4.24 statking/zephyr-7b-sft-full-orpo 1.616 FairMind/Llama-3-8B-4bit-UltraChat-Ita 0.296 bigcode/starcoderbase-7b 4.44 Severian/ANIMA-Phi-Neptune-Mistral-7B 1.692 NousResearch/Hermes-2-Theta-Llama-3-8B 0.330 deepseek-ai/deepseek-coder-6.7b-base 4.87 sethuiyer/Medichat-Llama3-8B 1.718
Table 1: Top 10 nearest neighbors among the 1,018 language models for each model listed in the first row. The values indicate the KL divergence measured in bits per byte (bpb), as defined in Section 3.5. These are computed using formula (3) in Section 2, by multiplying the original KL divergence by 0.001484. Tables for the nearest neighbors of the models labeled in the top panel of Fig. 1 are provided in Appendix I.
@@ 2.2 Log-likelihood vector
For a model $p_{i}$, the probability of generating a text $x$ is denoted $p_{i}(x)$. Following the convention in statistical model selection, we refer to $p_{i}(x)$ as the likelihood of model $p_{i}$ given the text $x$. The log-likelihood $\ell_{i}(x)=\log p_{i}(x)$ is then computed as
@@ 2.4 Kullback-Leibler divergence
The main results, namely (2) and (3), are proved in Appendix D using the theory of exponential family of distributions (Barndorff-Nielsen, 2014; Efron, 1978, 2022; Amari, 1982), similar to the discussion on the relationship between the norm of embeddings and KL divergence (Oyama et al., 2023). Although the concepts of model map and model coordinates have been discussed in statistics (Shimodaira, 1993; Shimodaira and Cao, 1998; Shimodaira, 2001), and there have been a few applications of model maps (Shimodaira and Hasegawa, 2005; Shimodaira and Terada, 2019), they have received little attention or use in practice.
@@ 2.5 Model coordinates
We primarily use $\bm{q}_{i}$ as the feature vector of model $p_{i}$ and refer to it as the model coordinates [^3] We refer to the Euclidean space where the vectors $\bm{\ell}_{i}$, $\bm{\xi}_{i}$, or $\bm{q}_{i}$, possibly scaled, reside as the log-likelihood space.. As shown in (3), the squared Euclidean distance in the $\bm{q}$-coordinate system approximates the KL divergence between language models [^4] That is, $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}\approx 2N\,\mathrm{KL}(p_{i},p_{j})$. For simplicity, we omit the constant scaling factor in expressions of this type., indicating that $\bm{q}_{i}$ represents the position of $p_{i}$ in the space of probability distributions. Since $\bm{\xi}_{i}$ differs from $\bm{q}_{i}$ only by an offset from the origin, $\bm{\xi}_{i}$ also serves as a model coordinate, and $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\|\bm{\xi}_{i}-\bm{\xi}_{j}\|^{2}$. However, we prefer $\bm{q}_{i}$ for its more interpretable components and thus adopt it throughout this paper.
@@ 3 Experimental Setup
Figure 4: Hierarchical clustering of the top 100 most-downloaded models, based on their feature vectors $\bm{q}_{i}$. Model names are color-coded by model type. KL divergence is reported in units of bits per byte (bpb).
@@ 3.1 Selection of text data
The texts used for computing the language models’ coordinates were extracted from the Pile Gao et al. (2020), with five categories of copyrighted material removed [^5] https://huggingface.co/datasets/monology/pile-uncopyrighted. This yielded a dataset $D$ consisting of 10,000 texts, each tagged with a category label from the Pile. Figure 2 visualizes these texts.
@@ 3.2 Selection of language models
We used $K=1{,}018$ language models in total. Of these, 1,000 were selected from models listed on Open LLM Leaderboard v1. Specifically, we considered CausalLM models ranging from 1B to 13B parameters and ranked them by their number of downloads over the 30 days preceding February 1, 2025666Hugging Face’s API provides the total number of downloads in the past 30 days.. We initially selected the top 1,100 models by download count and attempted log-likelihood calculations. Among these, 1,011 successfully produced valid log-likelihood values, and we chose the 1,000 most frequently downloaded from that set. In addition, we included 18 models from the DeepSeek language model series. We obtained information on model parameter sizes and architectures from the Leaderboard. Appendix G provides basic information on the selected models and details on how model types were defined. A complete list of models used in this study is given in Appendix L.
@@ 3.4 Obtaining the leaderboard scores
We obtained benchmark scores for the language models used in our experiments from Open LLM Leaderboard v1777https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard. Between April 2023 and June 2024, this leaderboard evaluated language models on six tasks: AI2 Reasoning Challenge (ARC) Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021a), TruthfulQA Lin et al. (2022a), Winogrande Sakaguchi et al. (2019), and GSM8K Cobbe et al. (2021). Along with individual task scores, we also use the average score across all six tasks, referred to as 6-TaskMean.
@@ 3.5 Byte-normalized KL divergence for cross-experiment comparison
Figure 6: Model maps color-coded by (Left) number of parameters and (Right) model creation date.
@@ 5.1 Benchmark scores and models
We use six benchmark scores from Open LLM Leaderboard v1, as described in Section 3.4. Our experiments are conducted on 996 models for which these benchmark scores are available111111 From the 1,018 models, we excluded four that lacked TruthfulQA scores and 18 from deepseek-ai, as their benchmark scores were incomplete due to the models being relatively new, leaving 996 for analysis. For simplicity, we denote the number of models as $K$. .
Pearson’s $r$ $0.946$ $0.909$ $0.932$ $0.901$ $0.941$ $0.884$ $0.953$ $0.989$
Spearman’s $\rho$ $0.948$ $0.956$ $0.934$ $0.884$ $0.948$ $0.857$ $0.960$ $0.974$
Table 2: Results of ridge regression for predicting benchmark scores from model coordinates. Predictions for 6-TaskMean and mean log-likelihood are also included. High correlation coefficients are observed across all settings. See Table 7 in Appendix K for the mean and standard deviation of correlation coefficients across the five data splits.
ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K 6-TaskMean
Pearson’s $r$ $0.453$ $0.598$ $0.346$ $0.072$ $0.508$ $0.239$ $0.395$
Spearman’s $\rho$ $0.432$ $0.467$ $0.422$ $0.048$ $0.512$ $0.364$ $0.400$
Table 3: Correlation between mean log-likelihoods and benchmark scores. The results show that perplexity has only moderate correlations on most tasks, especially compared to model coordinates in Table 2.
@@ 5.2 Setting for regression analysis
For each benchmark task, the dataset is given as $\{(\bm{q}_{1},v_{1}),\dots,(\bm{q}_{K},v_{K})\}$, where $\bm{q}_{i}\in\mathbb{R}^{N}$ is the double-centered log-likelihood vector of the language model $p_{i}$, and $v_{i}\in[0,100]$ is its corresponding benchmark score. We use ridge regression to predict each benchmark score. Let $\bm{Q}\in\mathbb{R}^{K\times N}$ be the matrix of explanatory variables, and let $\bm{v}=(v_{1},\ldots,v_{K})^{\top}\in\mathbb{R}^{K}$ be the vector for the target variable. The objective function with parameter $\bm{w}\in\mathbb{R}^{N}$ is given by:
$\mathcal{L}(\bm{w})=\|\bm{v}-\bm{Q}\bm{w}\|^{2}+\alpha\|\bm{w}\|^{2},$ (5)
@@
We partition the models into five folds based on model types and perform parameter training and benchmark score prediction [^12] We used the GroupKFold and RidgeCV implementations provided by scikit-learn Varoquaux et al. (2015).. To mitigate the effect of randomness, we repeat the data splitting with five different seeds and take the average of the predictions as the final predicted score. As evaluation metrics, we compute Pearson’s $r$ and Spearman’s $\rho$ to measure the correlation between the predicted and benchmark scores. Additionally, we conduct experiments by replacing the target variable with 6-TaskMean and mean log-likelihood, leading to a total of eight experimental settings. See Appendix K.1 for details.
5.3 Results and discussion
Regression analysis.
Table 2 summarizes the results of the regression analysis using $q$-coordinates. Across all benchmark tasks, both Pearson’s $r$ and Spearman’s $\rho$ are consistently high, indicating that the ridge regression model based on model coordinates achieves strong predictive performance. This trend holds not only for individual tasks but also for the 6-TaskMean, where the correlation coefficients remain high, as illustrated in the scatter plot in Fig. 8. Even when the target variable is the mean log-likelihood, defined as the simple average over the 10,000 texts, the regression model still achieves high correlation. This suggests that the model coordinates, despite being double-centered and excluding direct information about the mean, retain a sufficiently rich structure to accurately predict the average log-likelihood.
Figure 8: Scatter plot comparing predicted and actual 6-TaskMean scores for test sets (the dashed line indicates the identity line). Pearson’s $r$ is 0.953 and Spearman’s $\rho$ is 0.960, indicating strong predictive performance. Points are color-coded by the mean log-likelihood, with the color scale clipped at the 10th percentile. While higher mean log-likelihood values tend to correspond to higher benchmark scores, a few models with unusually high log-likelihoods due to data leakage (Section 4.2) deviate from this trend. Nevertheless, the predictions remain accurate overall. Scatter plots for individual benchmark tasks are shown in Fig. 18 in Appendix K.2.
Performance prediction based on perplexity.
The mean log-likelihood equals $-\log(\text{perplexity})$ and is often used as a performance indicator. We evaluated its correlation with benchmark scores, and Table 3 shows moderate correlations across all tasks, consistent with prior findings Liu et al. (2023a); Fang et al. (2025); Thrush et al. (2025). Unlike mean log-likelihood, which uniformly averages over texts, regression using the log-likelihood matrix $\bm{L}$ can assign task-dependent weights. As shown in Table 9 in Appendix K, its prediction accuracy is nearly the same as that of $\bm{Q}$. This indicates that combining $\bm{Q}$ and the mean in $\bm{L}$ offers little improvement. Accurate prediction with $\bm{Q}$ alone suggests that model positions already encode a structure aligned with benchmark performance.
@@ 6 Empirical Validation of Theory
In Section 2, we discussed model coordinates for text probability models. Appendix E extends this framework to token-sequence probability models. These theoretical results state that the squared Euclidean distance in the log-likelihood space approximates the KL divergence between models. To validate this, we first evaluate token-level conditional probability models (Section 6.1), since token-level experiments are generally easier to conduct than text-level experiments. We then extend our analysis to text probability models (Section 6.2). Additionally, in Section 6.3, we explore the relationship between model weight parameters and model coordinates.
@@ 6.3 Relationship between model weights and model coordinates
Figure 10: Visualization of 36 language models obtained by linearly interpolating pretrained model weights based on Llama-2-7b-hf. Each point is color-coded according to its mean log-likelihood. (Left) Models in the weight parameter space. (Right) Models in the log-likelihood space, represented by the $\bm{q}$-coordinate system.
For language models with the same architecture, comparison can also be conducted via weight parameters. We investigated how the structure of the log-likelihood space aligns with the structure of the weight-parameter space. We generated 36 new language models by linearly interpolating the weights of Llama-2-7b-hf, Llama-2-7b-chat-hf (Touvron et al., 2023b), and vicuna-7b-v1.5 (Zheng et al., 2023), using a $6\times 6$ grid of coefficients. For these 36 models, we computed text-generation log-likelihoods and visualized the resulting model coordinates in two dimensions using Principal Component Analysis (PCA). Figure 10 shows these 36 models in both weight space and log-likelihood space, where the two-dimensional grid structure is preserved. Interestingly, the mean log-likelihoods of the interpolated models closely follow the linear interpolation of the three base models’ mean log-likelihoods. This suggests that such interpolation can be a practical tool for exploring high-performing models. See Appendix J for details.